OpenStudy (anonymous):

Given equation A as 2x+1/4y=3 and equation B as 2/3x-y=6 , which expression will give the value of y?

OpenStudy (anonymous):

\[2x+\frac{ 1 }{ 4 }=3 \]

OpenStudy (anonymous):

\[ \frac{ 2 }{ 3 }x-y=6\]

OpenStudy (anonymous):

equation A: subtract 1/4 y to both sides, then divide by 2. What do you get?

OpenStudy (anonymous):

Thats where i get lost,i have no clue how to do this stuff

OpenStudy (anonymous):

\[2x + \frac{ 1 }{ 4 } y -\frac{ 1 }{ 4 }y = 3 -\frac{ 1 }{ 4 } y\] What's the answer?

OpenStudy (anonymous):

well, simplify it down :)

OpenStudy (anonymous):

ok,now how would i simplify that down? it looks easy but i dont get it

OpenStudy (anonymous):

what is 1/4 y - 1/4 y?

OpenStudy (anonymous):

is it just Y?

OpenStudy (anonymous):

hm no, for example, just assume that we know y = 4 so (1/4) *4 = 1, right? 1-1 = ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

yes, so, 1/4 y - 1/4 y = 0

OpenStudy (anonymous):

\[ 3-\frac{ y }{ 4 }\]

OpenStudy (anonymous):

No, hold on you're not done with equation A yet

OpenStudy (anonymous):

thats what i got when i did it all out but it looks wrong

OpenStudy (anonymous):

Ok, the old equation that we have was \[2x +\frac{ 1 }{ 4 y }= 3\] \[2x +\frac{ 1 }{ 4 }y -\frac{ 1 }{ 4 }y= 3 - \frac{ 1 }{ 4 } y\] \[2x + 0= 3 - \frac{ 1 }{ 4 } y\] Now you want x alone, so divide 2 to both sides.

OpenStudy (anonymous):

so x+0=1.5-1/2y?

OpenStudy (anonymous):

you can get rid of the 0 , it's extra space . and no \[1/5 - \frac{ 1 }{ 4} * \frac{ 1 }{ 2 }y\] Revise your answer?

OpenStudy (anonymous):

not the dash. 1.5, which is same as 3/2

OpenStudy (anonymous):

8 - 5y —— ? 40

OpenStudy (anonymous):

no, it would be \[x= 1.5 - \frac{ 1 }{ 8} y\] Now leave equation A like that for now. Simplify equation B

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 } x-y=6\]

OpenStudy (anonymous):

add y to both sides of equation B

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }x=6y\]

OpenStudy (anonymous):

the right side is wrong when you add 6 and y , it is the same as 6+y 6y means 6 * y so it's \[\frac{ 2 }{ 3 }x = 6+ y\] Do you know what you should do to get x alone?

OpenStudy (anonymous):

take 2/3 from both sides

OpenStudy (anonymous):

multiply 3 to both sides, then divide 2 to both sides

OpenStudy (anonymous):

x=9+y

OpenStudy (anonymous):

A − 3B A − B 2A − 3B A + B are my answers

OpenStudy (anonymous):

No, \[x = \frac{ 3 }{ 2} (6+y)\]

OpenStudy (anonymous):

then multiply 3/2 in for 6 and y, separately

OpenStudy (anonymous):

\[x = \frac{ 18 }{ 2 } y + \frac{ 3 }{ 2 }y\]

OpenStudy (triciaal):

Equation A 2x+1/4y=3 (3 -2 x)*4 = y = 12 - 8 x Equation B 2/3x-y=6 -y = 6 - (2/3 )x y = (2 x - 18)/3 question asked "which expression will give the value of y?" do you have a choice of expressions for the answer? It does not say solve for y when equation A = Equation B

OpenStudy (anonymous):

sorry no y iin the 18/2

OpenStudy (anonymous):

A − 3B A − B 2A − 3B A + B are the answer choices :( i have no clue what im doing

OpenStudy (anonymous):

i got b :/

OpenStudy (anonymous):

did you get the right answer or ...? I'm trying to find a way to explain it out

OpenStudy (anonymous):

i got it right..it was b

OpenStudy (anonymous):

so thank you

OpenStudy (anonymous):

it's actually A, when I solve everything out

OpenStudy (anonymous):

my way was correct, just that you did not need to find y

OpenStudy (anonymous):

@shadow.655

OpenStudy (anonymous):

Hope that helps D:

OpenStudy (anonymous):

ohh whoops,well its only one question.so thanks for your help